Two charges of equal magnitude exert an attractive force of 5.5×10-4 N on each other. If the magnitude of each charge is +1.8μ C, how far apart are the charges?
Answer #1

force = kq^2 / r^2

where q = charge

r= seperation

5.5*10^-4 = 9*10^9*(1.8*10^-6)^2 / r^2

r = 7.28 m

so the seperation between the charge = 7.28 m